OpenStudy (vannah):

8 (56)/6,plsz simplify that and show me your work plsz,medal is given

OpenStudy (anonymous):

4(56)/3

OpenStudy (vannah):

can you plsz show me how you got that

OpenStudy (anonymous):

you can solve it by using G.C.F

OpenStudy (chaise):

\[8 \frac{56}{6}\] Multiply the denominator (6) by the number out from (8), plus the numerator (56) then divide by the denominator (6) \[\frac{52}{3}\]

OpenStudy (chaise):

\[\frac{52}{3}=\frac{104}{6}\]